Please answer the following questions involving MIPS assembly code:

A) For the C statement below, what is the corresponding MIPS assembly code? Assume f, g are stored in S1 and S2.

f = g + (-f -5)

B)

For the C statement below, what is the corresponding MIPS assembly code? Assume i and j are assigned in registers S1 and S2 respectively and base of address of arrays A and B are in registers S6 and S7.

B[8] = A[i –j]

C)

For the instruction below, show the binary then hexadecimal representation of these instructions.

sw $t, 32 ($t2)

sll $s1, $2, 4

D)

For the lines below, what is the value of $t2 for the following sequence of instructions?

Sll $t2, $t0, 4

andi $t2, $t2, -1

Assume, t0 stores a value of x000020A1
Answer #1

A)

f = g + (-f -5)

addi $s1,$s1,5 # add 5 to f i.e, f+5 and assign it to $s1
sub $s1,$s2,$s1 # subtract f from g and assign it to f => f= g + (-f - 5) = g-(f+5)

B)

B[8] = A[i –j]

sub $s1,$s1,$s2 # subtract j from i and assign it to $s1 = i-j  
li $t1,8
li $t2,4

mul $t1,$t1,$t2 # $t1=32
mul $t3,$s1,$t2 # $t3 = (i-j)*4

add $t1,$t1,$s7 #add B array base address to $t1 which gives address B[8]
add $t3,$t3,$s6 #add A array address to $t3 which gives address A[i-j]

lw $t4,0($t3) #load A[i-j] to $t4
sw $t4,O($t1) # store $t4 in address $t1 i.e B[8]

D) value of $t2= 0x000

sll $t2, $t0, 4 #shift left 4 bits bu the number is hex so it will shift to => t2 = 00020A10
andi $t2, $t2, -1 # do and operation $t2,-1 =>t2= 0x000